OpenStudy (anonymous):

What is thee area of TUVW, if TU =7 WV =15 and UV = 10 I know the answer is 66, but not sure how height was calculated to be 6

OpenStudy (campbell_st):

what type of quadrilateral

OpenStudy (anonymous):



sam (.sam.):

\[\huge A=\frac{1}{2}(WV+TU)(TW)\]

OpenStudy (anonymous):

got that - my question is how is tw calculated to be 6

OpenStudy (campbell_st):

A = 1/2 x height x ( sum of parallel sides)

sam (.sam.):

\[\huge TW=\frac{2A}{WV+TU}\]

OpenStudy (anonymous):

No, I understand that - wwhat I want to see is calculations. I am not sure of how my height of this figure was 6 son's sat prep and i know that thee answer was a multiple of11 based on other numbers so thereofre - 66 had to be the area

OpenStudy (campbell_st):

use pythagoras draw a perpendicular from U to WV... call the point Z then (TW)^2=10^2 - 8^2 TW = 6

OpenStudy (campbell_st):

TW and UZ have the same height.

OpenStudy (anonymous):

ah ha :) I drew that darn line and like an idiot wrote 7 - head was thinking square. Great explanation Campbell - really appreciate that.